# Dynamics of Pseudoentanglement

### v1

@@ VI.2 Outlook
Our work opens a number of directions for future research. First of all, it would be desirable to improve and/or rigorize our results in several ways: by tightening bounds, especially on the mixing times in the local circuit model and on pseudothermalization without random phases; by proving or disproving our Conjecture 1 on the relaxation times in the local circuit model; and by proving the existence of cutoff phenomena in our Markov chains, for which we only showed a necessary condition (divergent ratio of mixing time and relaxation time). These improvements could have significant qualitative consequences on the physics. For example, one may hope to show that, in the local circuit model for pseudoentanglement spreading, Sec. IV.3, the $m$-copy pseudothermalization transitions are sharply separated from one another, in the sense that the difference of consecutive mixing times $t_{\rm mix}^{(m+1)}-t_{\rm mix}^{(m)}$ is sufficiently larger than the relaxation time $t_{\rm rel}^{(m)}$.
@@
Another direction for future work is to identify physically realistic models of dynamics where pseudoentanglement is achieved in polynomial time. In the all-to-all model of Sec. IV.2 that time scale is exponential for trivial reasons; in the local model of Sec. IV.3, the $m$-copy pseudothermalization time is polynomial in system size, but grows at least linearly with $m$, so that convergence to a pseudoentangled ensemble, indistinguishabile from Haar-random with any polynomial number of copies $m=O({\rm poly}(N))$, requires super-polynomial time. This can be a mild super-polynomial scaling, e.g. $N^{\log N}$. Nonetheless it would be interesting to identify models of dynamics that generate pseudoentanglement in polynomial time. Such models cannot be random local automaton circuits, due to the counting bound, Eq. (30). Pseudorandom subset(-phase) state ensembles can be prepared efficiently by using quantum-secure pseudorandom permutations and phase functions Aaronson et al. (2023). In this work we have opted for truly random subsets, as these make for simpler models of circuit dynamics and are more in keeping with the spirit of quantum thermalization; it would be interesting to identify pseudorandom models of dynamics that can produce pseudoentanglement in polynomial time while maintaining a simple, “random-looking”, physically reasonable structure. A related construction of pseudorandom unitaries was recently shown to not be robust to perturbations Haug et al. (2023), which might suggest intrinsic difficulties for this direction.
@@
More broadly, subset(-phase) states are only one class of pseudoentangled state ensembles. Due to their structure, these ensembles suggest automaton circuits as a natural models of dynamics. It would be interesting to identify different ensembles of pseudoentangled states and the types of quantum dynamics that might naturally produce them.
@@
Recent works have introduced related notions of ‘pseudocoherence’ Haug et al. (2023), ‘pseudomagic’ Gu et al. (2023), etc, which are analogous to pseudoentanglement but based on different resources (coherence, magic, etc). A natural extensions of our work would be to study the dynamics of other such pseudoresources in appropriate circuit models.
@@
It is also interesting to ask whether, in analogy with measurement-induced entanglement phase transitions Li et al. (2018); Skinner et al. (2019); Gullans and Huse (2020); Iaconis et al. (2020), one might obtain phase transitions in the dynamics of pseudoentanglement by enriching our automaton circuits with suitable projective measurements and other elements. For example, while the size of subsets is invariant under automaton gates, it generally decreases under $Z$ measurements, and can increase under Hadamard gates and $X$ measurements. Can the competition of these elements give rise to nontrivial phase transitions in the pseudoentanglement properties of late-time states?
@@
Finally, the possible emergence of cutoff phenomena in pseudothermalization raises more general questions about the sharpness of thermalization in generic quantum dynamics. Are sharp transitions in distinguishability a generic feature of chaotic quantum dynamics Kastoryano et al. (2012); Oh and Kais (2023)? What are the right tools to analyze them? Can there particularly structured classes of time evolutions (e.g., with symmetry) exhibit different levels of sharpness? We leave these questions as an interesting target for future research.
@@ Acknowledgements.
We thank Nick Hunter-Jones, Ethan Lake, and Scott Aaronson for stimulating discussions. We are especially grateful to Nick Hunter-Jones for pointing us to Ref. Gowers (1996). Some numerical simulations were carried out on HPC resources provided by the Texas Advanced Computing Center (TACC) at the University of Texas at Austin.
@@ Appendix A Proof of Proposition 1
$\left\|\rho^{(m)}_{\mathcal{E}_{p}}-\rho^{(m)}_{\rm Haar}\right\|_{\rm tr}$ (45)
@@ A.1 Preliminaries
Eq. (45) takes place within the symmetric sector of the $m$-fold replicated Hilbert space $\mathcal{H}^{\otimes m}$. It is thus helpful to introduce an orthonormal basis of that subspace:
@@
where $\mathbf{z}\in[D]^{m}$ is an $m$-tuple of basis states and ${\rm type}(\mathbf{z})\in[m+1]^{D}$ is a vector that counts the multiplicity of each basis state $z$ in the $m$-tuple $\mathbf{z}$ (the notation is borrowed from Ref. Aaronson et al. (2023)): i.e., if ${\rm type}(\mathbf{z})=\mathbf{T}$, then $T_{i}$ is the number of times the basis state $i\in[D]$ appears in $\mathbf{z}$. As a consequence we have $\sum_{i}T_{i}=m$. Finally, we used a shorthand for the multinomial coefficient
$\binom{m}{\mathbf{T}}=\frac{m!}{T_{0}!\cdots T_{D-1}!}.$
@@ A.2 Proof
Our proof builds on Ref. Aaronson et al. (2023), specifically the two facts below:
@@
$\left\|\rho^{(m)}_{\mathcal{E}_{p}}-\rho^{(m)}_{\rm Haar}\right\|_{\rm tr}$ $=\left\|\sum_{S\in\Sigma_{K}}(p(S)-\pi_{K}(S))\rho_{S,{\rm unique}}^{(m)}\right\|_{\rm tr}$
$\qquad+O\left(\frac{m^{2}}{K}\right).$ (50)
Now, using the definition of unique-type state Eq. (49) and orthonormality of the unique-type states, we arrive at
@@ Appendix B Properties of the $\Phi$ maps
$\Phi_{c\leftarrow b}\circ\Phi_{b\leftarrow a}[p](S^{\prime\prime})$ $=\binom{a}{b}^{-1}\binom{b}{c}^{-1}\sum_{\begin{subarray}{c}S,S^{\prime}:\\ S^{\prime\prime}\subset S^{\prime}\subset S\end{subarray}}p(S)$ (58)
$=\binom{a}{b}^{-1}\binom{b}{c}^{-1}\binom{a-c}{b-c}\binom{a}{c}$
×$\qquad\times\Phi_{c\leftarrow a}[p](S^{\prime\prime}),$ (59)
@@
$=\binom{K}{m}^{-1}\sum_{S:\,S^{\prime}\subset S}\sum_{\begin{subarray}{c}\sigma\in\mathcal{S}_{D}:\\ \sigma(S_{0})=S\end{subarray}}\mathbb{P}_{t}(\sigma)$ (64)
@@
$\Phi_{m\leftarrow K}\left[p_{t,S_{0}}^{(K)}\right](S^{\prime})$ $=\binom{K}{m}^{-1}\sum_{S_{0}^{\prime}\subset S_{0}}\sum_{\begin{subarray}{c}\sigma\in\mathcal{S}_{D}:\\ S^{\prime}=\sigma(S_{0}^{\prime})\end{subarray}}\mathbb{P}_{t}(\sigma)$ (70)
@@
where we have recognized the definition of $p_{t,S_{0}^{\prime}}^{(m)}(S^{\prime})$, the Markov chain induced on $\Sigma_{m}$, with initial state $S_{0}^{\prime}$ averaged uniformly over all subsets of $S_{0}$ of the appopriate cardinality, as in Eq. (61). ∎
@@
$\Delta_{m}\equiv\|\Phi_{m\leftarrow K}[p]-\pi_{m}\|_{\rm tv}$ (72)
@@
$\Delta_{m-1}$ $=\|\Phi_{m-1\leftarrow K}[p-\pi_{K}]\|_{\rm tv}$
$=\|\Phi_{m-1\leftarrow m}\circ\Phi_{m\leftarrow K}[p-\pi_{K}]\|_{\rm tv}$
$\leq\|\Phi_{m\leftarrow K}[p-\pi_{K}]\|_{\rm tv}=\Delta_{m}.$ (73)
The first line uses the fact that $\Phi_{m\leftarrow K}[\pi_{K}]=\pi_{m}$ for all $m<K$; the second uses the semigroup law of $\Phi$ maps, Fact B.2; the last one uses monotonicity of total variation norm under stochastic maps: $\|\Phi[\cdots]\|_{\rm tv}\leq\|\cdots\|_{\rm tv}$. ∎
@@ Appendix C Proof of Proposition 2
We begin by recalling a useful fact proven in Ref. Giurgica-Tiron and Bouland (2023):
@@
See Theorem 1 in Ref. Giurgica-Tiron and Bouland (2023). ∎
@@
$\left\|\sum_{S}[p(S)-\pi_{K}(S)]\outerproduct{\psi_{S}}{\psi_{S}}^{\otimes m}\right\|_{{\rm tr}}$ (75)
@@
where |$\ket{S^{\prime}}$ represents the unique-type basis state |$\ket{\mathbf{T}}$ with 1’s only on states in $S^{\prime}$. As also noted in Appendix A, in the limit of $D\gg m$ most types are unique. This is helpful as it allows us to replace the state |$\ket{\psi_{S}}$ in Eq. (75) by its unique-type counterpart, |$\ket{\psi_{S,{\rm uni}}}$, with a small error. To show this, we first need a simple fact about the trace distance between pure states:
@@
See Fact 2 in Ref. Aaronson et al. (2023) (set $\rho_{1}=\outerproduct{\phi}{\phi}$ |, $\rho_{2}=\outerproduct{\chi}{\chi}$ |). ∎
@@
$\innerproduct{\psi_{S}^{\otimes m}}{\psi_{S,{\rm uni}}^{(m)}}$ $=K^{-m/2}\binom{K}{m}^{-1/2}$
×$\qquad\times\sum_{\mathbf{T}\in\mathcal{T}_{S}^{m}}\binom{m}{\mathbf{T}}^{1/2}\delta(\mathbf{T}\text{ is unique})$
@@
Using some straightforward algebra and the ‘birthday asymptotics’ $\frac{K!}{(K-m)!}=K^{m}\left(1+O(m^{2}/K)\right)$ (see e.g. Fact 1 in Ref. Giurgica-Tiron and Bouland (2023)) we finally get
@@
$\left\|\sum_{S\in\Sigma_{K}}(p(S)-\pi_{K}(S))\outerproduct{\psi_{S,{\rm uni}}^{(m)}}{\psi_{S,{\rm uni}}^{(m)}}\right\|_{\rm tr}$ (82)
now up to error $O(m/\sqrt{K})$. Finally, plugging in the definition of unique-type pure states, Eq. (77), yields
$\left\|\sum_{S\in\Sigma_{K}}(p(S)-\pi_{K}(S))\binom{K}{m}^{-1}\sum_{\begin{subarray}{c}S^{\prime},S^{\prime\prime}\in\Sigma_{m}:\\ S^{\prime},S^{\prime\prime}\subset S\end{subarray}}\outerproduct{S^{\prime}}{S^{\prime\prime}}\right\|_{\rm tr}$ (85)
@@
$\binom{K}{m}^{-1}\binom{K}{m+\delta}\Phi_{m+\delta\leftarrow K}[p-\pi](S^{\prime}\cup S^{\prime\prime})$ (89)

### v3

@@ 6.3 Outlook
Our work opens a number of directions for future research. First of all, it would be desirable to improve and/or rigorize our results in several ways: by tightening bounds, especially on the mixing times in the local circuit model and on pseudoentanglement without random phases; by proving or disproving our Conjecture 1 on the relaxation times in the local circuit model; and by proving the existence of cutoff phenomena in our Markov chains, for which we only showed a necessary condition (divergent ratio of mixing time to relaxation time). These improvements could have significant qualitative consequences. For example, one may hope to show that, in the local circuit model for pseudoentanglement spreading, Sec. 4.3, the $m$-copy distinguishability transitions are sharply separated from one another, in the sense that the difference of consecutive mixing times $t_{\rm mix}^{(m+1)}-t_{\rm mix}^{(m)}$ is sufficiently larger than the relaxation time $t_{\rm rel}^{(m)}$. This would give a cascade of distinguishability transitions, one for each number of copies $m$ that the observer can access, sharply separated in time from each other.
@@
Another direction for future work is to identify physically realistic models of dynamics where pseudoentanglement is achieved in polynomial time. In the all-to-all model of Sec. 4.2 that time scale is exponential for trivial reasons; in the local model of Sec. 4.3, the $m$-copy indistinguishability time is polynomial in system size, but grows at least linearly with $m$, so that convergence to a pseudoentangled ensemble, indistinguishable from Haar-random with any polynomial number of copies $m=O({\rm poly}(N))$, requires super-polynomial time. This is a consequence of demanding information-theoretic pseudorandomness, i.e. vanishing trace distance [Eq. (3)], along with a gate set of polynomial size (this follows from the counting bound, Fact 4.1). It would be interesting to identify physically reasonable models of dynamics that generate pseudoentanglement in polynomial time, under the weaker condition of computational pseudorandomness [Eq. (2)]. Indeed pseudorandom subset(-phase) state ensembles can be prepared efficiently by using quantum-secure pseudorandom permutations and phase functions aaronson_quantum_2023 . In this work we have considered truly random circuits, as these make for simpler models of dynamics and are more in keeping with the spirit of quantum thermalization; however it would be interesting to identify pseudorandom models of dynamics that can produce pseudoentanglement in polynomial time while maintaining a simple, local, “random-looking” structure. A related construction of pseudorandom unitaries was recently shown to not be robust to perturbations haug_pseudorandom_2023 , which might present an obstruction for this direction.
@@
More broadly, subset(-phase) states are only one class of pseudoentangled state ensembles. Due to their structure, these ensembles suggest automaton circuits as a natural models of dynamics. It would be interesting, given different ensembles of pseudoentangled states (for instance a recent construction based on shallow circuits of pseudorandom unitaries acting on large blocks of qubits schuster_random_2024 ), to identify models of quantum dynamics that might naturally produce them as late-time steady states.
@@
Recent works have introduced related notions of ‘pseudocoherence’ haug_pseudorandom_2023 , ‘pseudomagic’ gu_little_2023 , etc, which are analogous to pseudoentanglement but based on different resources (coherence, magic, etc). A natural extensions of our work would be to study the dynamics of other such pseudoresources in appropriate circuit models.
@@
It is also interesting to ask whether, in analogy with measurement-induced entanglement phase transitions li_quantum_2018 ; skinner_measurement-induced_2019 ; gullans_dynamical_2020 ; iaconis_measurement-induced_2020 , one might obtain phase transitions in the dynamics of pseudoentanglement by enriching our automaton circuits with suitable projective measurements and other elements. For example, while the size of subsets is invariant under automaton gates, it generally decreases under $Z$ measurements, and can increase under Hadamard gates and $X$ measurements. Can the competition of these elements give rise to nontrivial phase transitions in the pseudoentanglement properties of late-time states? Such transitions might carry implications for the hardness of detecting general measurement-induced phenomena, as discussed earlier.
@@
Finally, the possible emergence of cutoff phenomena in ensemble pseudothermalization raises more general questions about the sharpness of thermalization in generic quantum dynamics. Are sharp transitions in distinguishability a generic feature of chaotic quantum dynamics kastoryano_cutoff_2012 ; oh_cutoff_2023 ; balasubramanian_glassy_2024 ? What are the right tools to analyze them? Can particularly structured classes of time evolutions (e.g., with symmetry) exhibit different levels of sharpness? We leave these questions as an interesting target for future research.
Note added. Shortly after the first version of this manuscript appeared on arXiv, two new results on the spectral gaps of random reversible classical circuits have appeared. Ref. he_pseudorandom_2024 studied a model of geometrically local 3-bit gates very similar to ours and proved (in our notation) an upper bound on the relaxation time $t_{\rm rel}\leq\tilde{O}(mN)$. Subsequently, Ref. chen_incompressibility_2024 considered 3-local, but geometrically non-local, random reversible circuits and showed an upper bound $t_{\rm rel}\leq\tilde{O}(N)$ for all $m=O({\rm poly}(N))$. They conjecture that the bounds should apply to the geometrically-local setting as well, which would prove our Conjecture 1.
@@ Acknowledgements.
We thank Scott Aaronson, Tudor Giurgica-Tiron, Nick Hunter-Jones, Ethan Lake, and Shivan Mittal for stimulating discussions. We are especially grateful to Nick Hunter-Jones for pointing us to Ref. gowers_almost_1996 and to Shivan Mittal for sharing his results mittal_private_2024 on the gap of a Hamiltonian discussed in Appendix D.3. Some numerical simulations were carried out on HPC resources provided by the Texas Advanced Computing Center (TACC) at the University of Texas at Austin.
@@ Appendix A Proof of Proposition 1
$D_{\rm tr}\left(\rho^{(m)}_{\mathcal{E}_{p}},\rho^{(m)}_{\rm Haar}\right)=\frac{1}{2}\left\|\rho^{(m)}_{\mathcal{E}_{p}}-\rho^{(m)}_{\rm Haar}\right\|_{\rm tr}$ (46)
@@ A.1 Preliminaries
Eq. (46) involves states in the symmetric sector of the $m$-fold replicated Hilbert space $\mathcal{H}^{\otimes m}$. It is thus helpful to introduce an orthonormal basis of that subspace:
@@
where $\mathbf{z}\in[D]^{m}$ is an $m$-tuple of basis states and ${\rm type}(\mathbf{z})\in[m+1]^{D}$ is a vector that counts the multiplicity of each basis state $z$ in the $m$-tuple $\mathbf{z}$ (the notation is borrowed from Ref. aaronson_quantum_2023 ): i.e., if ${\rm type}(\mathbf{z})=\mathbf{T}$, then $T_{z}$ is the number of times the basis state $z\in[D]$ appears in $\mathbf{z}$. As a consequence we have $\sum_{z=0}^{D-1}T_{z}=m$. Finally, we used a shorthand for the multinomial coefficient
$\binom{m}{\mathbf{T}}=\binom{m}{T_{0},\dots T_{D-1}}=\frac{m!}{T_{0}!\cdots T_{D-1}!}.$
@@ A.2 Proof
Our proof builds on Ref. aaronson_quantum_2023 , specifically the two facts below:
@@
$\left\|\rho^{(m)}_{\mathcal{E}_{p}}-\rho^{(m)}_{\rm Haar}\right\|_{\rm tr}$ $=\left\|\sum_{S\in\Sigma_{K}}(p(S)-\pi_{K}(S))\rho_{S,{\rm unique}}^{(m)}\right\|_{\rm tr}+O\left(\frac{m^{2}}{K}\right).$ (51)
Now, using the definition of unique-type state Eq. (50) and orthonormality of the {|$\{\ket{S^{\prime}}\}_{S^{\prime}\in\Sigma_{m}}$ states, we arrive at
@@ Appendix B Properties of the $\Phi$ maps
$\Phi_{c\leftarrow b}\circ\Phi_{b\leftarrow a}[p](S^{\prime\prime})$ $=\binom{a}{b}^{-1}\binom{b}{c}^{-1}\sum_{\begin{subarray}{c}S,S^{\prime}:\\ S^{\prime\prime}\subset S^{\prime}\subset S\end{subarray}}p(S)$
$=\binom{a}{b}^{-1}\binom{b}{c}^{-1}\binom{a-c}{b-c}\binom{a}{c}\times\Phi_{c\leftarrow a}[p](S^{\prime\prime}),$ (55)
@@
$=\binom{K}{m}^{-1}\sum_{S:\,S^{\prime}\subset S}\sum_{\begin{subarray}{c}\sigma\in\mathcal{S}_{D}:\\ \sigma(S_{0})=S\end{subarray}}\mathbb{P}_{t}(\sigma)$
@@
$\Phi_{m\leftarrow K}\left[p_{t,S_{0}}^{(K)}\right](S^{\prime})$ $=\binom{K}{m}^{-1}\sum_{S_{0}^{\prime}\subset S_{0}}\sum_{\begin{subarray}{c}\sigma\in\mathcal{S}_{D}:\\ S^{\prime}=\sigma(S_{0}^{\prime})\end{subarray}}\mathbb{P}_{t}(\sigma)$
@@
where we have recognized the definition of $p_{t,S_{0}^{\prime}}^{(m)}(S^{\prime})$, the Markov chain induced on $\Sigma_{m}$, with initial state $S_{0}^{\prime}$. We see that $S_{0}^{\prime}$ is averaged uniformly over all subsets of $S_{0}$ of the appropriate cardinality, as in Eq. (57). ∎
@@
$\Delta_{m}\equiv D_{\rm tv}(\Phi_{m\leftarrow K}[p],\pi_{m})$ (60)
@@
$\Delta_{m-1}$ $=\frac{1}{2}\|\Phi_{m-1\leftarrow K}[p-\pi_{K}]\|_{\rm tv}$
$=\frac{1}{2}\|\Phi_{m-1\leftarrow m}\circ\Phi_{m\leftarrow K}[p-\pi_{K}]\|_{\rm tv}$
$\leq\frac{1}{2}\|\Phi_{m\leftarrow K}[p-\pi_{K}]\|_{\rm tv}=\Delta_{m}.$ (61)
The first line uses the fact that $\Phi_{m\leftarrow K}[\pi_{K}]=\pi_{m}$ for all $m<K$; the second uses the semigroup law of $\Phi$ maps, Fact B.2; the last one uses monotonicity of total variation distance under stochastic maps: $D_{\rm tv}(\Phi(p),\Phi(q))\leq D_{\rm tv}(p,q)$. ∎
@@ Appendix C Proof of Proposition 2
We begin by recalling a useful fact proven in Ref. giurgica-tiron_pseudorandomness_2023 :
@@
See Theorem 1 in Ref. giurgica-tiron_pseudorandomness_2023 . ∎
@@
$\frac{1}{2}\left\|\sum_{S}[p(S)-\pi_{K}(S)]\outerproduct{\psi_{S}}{\psi_{S}}^{\otimes m}\right\|_{{\rm tr}}$ (63)
@@
where |$\ket{S^{\prime}}$ stands for the type state |$\ket{\mathbf{T}}$ where $\mathbf{T}$ is the unique type whose 1’s are on the states in $S^{\prime}$ ($T_{z}=1$ iff $z\in S^{\prime}$, $T_{z}=0$ otherwise). As also noted in Appendix A, in the limit of $D\gg m$ most types are unique. This is helpful as it allows us to replace the state |$\ket{\psi_{S}}$ in Eq. (63) by its unique-type counterpart, |$\ket{\psi_{S,{\rm uni}}}$, with a small error. To show this, we first need a simple fact about the trace distance between pure states:
@@
See Fact 2 in Ref. aaronson_quantum_2023 (set $\rho_{1}=\outerproduct{\phi}{\phi}$ |, $\rho_{2}=\outerproduct{\chi}{\chi}$ |). ∎
@@
$\innerproduct{\psi_{S}^{\otimes m}}{\psi_{S,{\rm uni}}^{(m)}}$ $=K^{-m/2}\binom{K}{m}^{-1/2}\sum_{\mathbf{T}\in\mathcal{T}_{S}^{m}}\binom{m}{\mathbf{T}}^{1/2}\delta(\mathbf{T}\text{ is unique})$
@@
Using some straightforward algebra and the ‘birthday asymptotics’ $\frac{K!}{(K-m)!}=K^{m}\left(1+O(m^{2}/K)\right)$ (see e.g. Fact 1 in Ref. giurgica-tiron_pseudorandomness_2023 ) we finally get
@@
$\frac{1}{2}\left\|\sum_{S\in\Sigma_{K}}(p(S)-\pi_{K}(S))\outerproduct{\psi_{S,{\rm uni}}^{(m)}}{\psi_{S,{\rm uni}}^{(m)}}\right\|_{\rm tr}$ (70)
now up to error $O(m/\sqrt{K})$. Finally, plugging in the definition of unique-type pure states, Eq. (65), yields (up to the $1/2$ prefactor)
$\left\|\sum_{S\in\Sigma_{K}}(p(S)-\pi_{K}(S))\binom{K}{m}^{-1}\sum_{\begin{subarray}{c}S^{\prime},S^{\prime\prime}\in\Sigma_{m}:\\ S^{\prime},S^{\prime\prime}\subset S\end{subarray}}\outerproduct{S^{\prime}}{S^{\prime\prime}}\right\|_{\rm tr}$
@@
$\binom{K}{m}^{-1}\binom{K}{m+\delta}\Phi_{m+\delta\leftarrow K}[p-\pi_{K}](S^{\prime}\cup S^{\prime\prime})$ (72)